OpenStudy (anonymous):

Hey... Translation rules? Write the translation rule that maps P(-4,2) onto P'(-1,1)

OpenStudy (mathmath333):

Translation Rule \(\large \color{black}{\begin{align}-4+a&=-1\hspace{.33em}\\~\\ 2+b&=1\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}(x+a,y+b)\implies (x',y')\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

(x+3, y+-1) ?

OpenStudy (anonymous):

Negative.

OpenStudy (mathmath333):

do u have options ?

OpenStudy (anonymous):

i don't have options.

OpenStudy (mathmath333):

then it should be \(\large \color{black}{\begin{align}(x,y)\implies (x+3,y-1)\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

Oh okay! Thanks!